A random sample is drawn from a population of known standard deviation 11.3. Construct a 90% confidence interval for the population mean based on the information given (not all of the information given need be used).


n = 36, x ̅=105.2, s = 11.2


n = 100, x ̅=105.2, s = 11.2


1
Expert's answer
2021-03-01T07:10:42-0500

a.

"90\%CI=(\bar x-1.645\frac{s}{\sqrt{n}},\;\bar x+1.645\frac{s}{\sqrt{n}})"


"=(105.2-1.645\frac{11.2}{\sqrt{36}},\;105.2+1.645\frac{11.2}{\sqrt{36}})"


"=(102.13,\;108.27)."



b.

"90\%CI=(\bar x-1.645\frac{s}{\sqrt{n}},\;\bar x+1.645\frac{s}{\sqrt{n}})"


"=(105.2-1.645\frac{11.2}{\sqrt{100}},\;105.2+1.645\frac{11.2}{\sqrt{100}})"


"=(103.36,\;107.04)."